2. An object of mass 10 kg is whirled round a horizontal circle of radius 4 m by a revolving string inclined to the vertical. If the uniform speed of the object is 5 m/s, calculate:



(i) the tension in the string



(ii) the angle of inclination of the string to the vertical.

1
Expert's answer
2022-01-17T16:20:29-0500

"\text{forces acting on an object:}" "\text {}"

"F_g-\text{gravity}"

"F_g=mg=9.8\u00d710=98N"

"F_c-\text{force centripetal acceleration}"

"F_c=ma;a=\frac{v^2}{R}=\frac{25}{4}=6.25m/s^2"

"F_c=10\u00d76.25=62.5N"

"T-\text{string tension}"

"\text{X-axis projection}"

"T_x+F_g=0"

"|Tx|=F_g=98N"

"\text{Y-axis projection}"

"T_y= F_c"

"T_y=62.5N"

"T=\sqrt{T^2_x+T^2_y}=\sqrt{62.5^2+98^2}\approx116.23N"

"T=\sqrt{T^2_x+T^2_y}=\sqrt{62.5^2+98^2}\approx116.23N""T=\sqrt{T^2_x+T^2_y}=\sqrt{62.5^2+98^2}\approx116.23N"

"\alpha-\text{angle of inclination of the string to the vertical}"

"\tan{\alpha}=\frac{F_c}{T_x}=\frac{62.5}{98}\approx0.64"

"\alpha=32.6\degree"

Answer:"T=116.23N;\alpha=32.6\degree"